# Integral canonical models of exceptional Shimura varieties

### v1

@@ 2.3 Natural boundary and local systems
For $\mathcal{S}/\mathcal{O}$ smooth over a Dedekind domain and $\bar{\mathcal{S}}$ a log smooth compactification, both the naturality of the boundary and the admissibility condition on $\mathcal{O}$-morphisms to $\mathcal{S}$ can be checked using an $\ell$-adic local system $\mathbb{V}_{\ell}$ on $\mathcal{S}$. The key property such a local system must have is “nonextendability” over the generic fiber, meaning it does not extend to any nontrivial partial compactification. This notion is treated carefully in [brunebarbe, §3] (where it is referred to as “maximality”); the following definition is easily seen to be equivalent to the definition therein.
@@
Let $Y/F$ be a normal variety with an $\ell$-adic local system $\mathbb{V}_{\ell}$. We say $\mathbb{V}_{\ell}$ has faithful geometric local monodromy if for some (hence any) resolution $\pi:S\to Y$ and log smooth compactification $S\subset\bar{S}$, the geometric inertia at the generic point of every stratum of the boundary $D$ acts faithfully on $\pi^{*}\mathbb{V}_{\ell}$. We say an $\ell$-adic local system on an integral model has faithful geometric local monodromy if the restriction to the generic fiber does.
Note that faithfulness of geometric local monodromy can be checked over the algebraic closure $\bar{F}$, or even over $\mathbb{C}$ for a choice of embedding $F\subset\mathbb{C}$. Faithfulness of geometric local monodromy restricts well:
@@
    $\mathbb{V}_{\ell}$ has faithful geometric local monodromy.
@@
    For any curve $C/\mathbb{C}$ and any proper morphism $f:C\to Y$, $f^{*}\mathbb{V}_{\ell}$ has faithful geometric local monodromy.
@@
    For any normal variety $X/\mathbb{C}$ and any proper morphism $f:X\to Y$, $f^{*}\mathbb{V}_{\ell}$ has faithful geometric local monodromy.
The next lemma interpets naturality of the boundary and admissibility in terms of the local system as promised.
@@
Suppose $\mathcal{Y}/\mathcal{O}$ is a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$, and $\mathbb{V}_{\ell}$ an $\ell$-adic local system on $\mathcal{Y}$ such that $\mathbb{V}_{\ell}|_{Y}$ has faithful geometric local monodromy. Then
@@
    $\bar{\mathcal{S}}$ has natural boundary.
@@
    For $\mathcal{X}/\mathcal{O}$ regular, a morphism $f_{K}:X\to Y$ over the generic point which meets the locus of $Y$ where $\pi$ is an isomorphism is $\mathcal{O}$-admissible if and only if $f_{K}^{*}\mathbb{V}_{\ell}$ extends to $\mathcal{X}$.
@@
For 1, the geometric inertia of every stratum acts faithfully on $\pi_{K}^{*}\mathbb{V}_{\ell}$, and this action factors through $\pi_{1}^{{\mathrm{et}}}(\mathcal{S},\bar{s})\to\pi_{1}^{{\mathrm{et}}}(\mathcal{Y},\bar{y})$, so it follows that the boundary is natural.
@@ 2.4 Integral canonical models of smooth varieties
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ and $\mathcal{Y}/\mathcal{O}$ a quasiprojective smooth $\mathcal{O}$-scheme. We say $\mathcal{Y}/\mathcal{O}$ is an integral canonical model if for every closed point $v$ of $\operatorname{Spec}\mathcal{O}$, the base-change $\mathcal{Y}_{v}/\mathcal{O}_{v}$ satisfies the extension property over $\mathcal{O}_{v}$.
@@
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$. Let $\mathcal{S},\mathcal{S}^{\prime}/\mathcal{O}$ be two smooth integral canonical models and $f_{F}:S\to S^{\prime}$ a morphism over the generic point.
@@
    Suppose further that $\mathcal{S}^{\prime}$ admits a log smooth compactification and an $\ell$-adic local system $\mathbb{V}_{\ell}$ such that $\mathbb{V}_{\ell}|_{S^{\prime}}$ has faithful geometric local monodromy. Then $f_{F}$ extends to a morphism $f:\mathcal{S}\to\mathcal{S}^{\prime}$ if and only if $f_{F}^{*}(\mathbb{V}_{\ell})$ extends to $\mathcal{S}$.
@@ 2.5 Integral canonical models of normal varieties
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$. Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme.
@@
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$, $Y/F$ a normal variety, $\pi_{F}:S\to Y$ a resolution, and $\pi:\mathcal{S}\to\mathcal{Y}$ a model of $\pi_{F}$ defined over $\mathcal{O}$. Then $\pi$ is a uniform resolution over a (nonempty) open set of $\mathcal{O}$.
@@
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ and $\mathcal{Y}/\mathcal{O}$ a normal quasiprojective flat $\mathcal{O}$-scheme. We say $\mathcal{Y}/\mathcal{O}$ is an integral canonical model if:
@@
    Suppose further that $\mathcal{Y}^{\prime}$ admits a uniform resolution $\pi^{\prime}:\mathcal{S}^{\prime}\to\mathcal{Y}^{\prime}$, a log smooth compactification of $\mathcal{S}^{\prime}$, and that $\mathcal{Y}^{\prime}$ has an $\ell$-adic local system $\mathbb{V}_{\ell}$ such that $\mathbb{V}_{\ell}|_{Y^{\prime}}$ has faithful geometric local monodromy. Then $f_{F}$ extends to a morphism $f:\mathcal{Y}\to\mathcal{Y}^{\prime}$ if and only if $f_{F}^{*}(\mathbb{V}_{\ell})$ extends to $\mathcal{Y}$.
@@
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$ Let $\mathcal{Y},\mathcal{Y}^{\prime}/\mathcal{O}$ be normal finite type flat $\mathcal{O}$-schemes, and $\pi:\mathcal{S}\to\mathcal{Y}$ a uniform resolution. Assume $\mathcal{Y}^{\prime}/\mathcal{O}$ is quasiprojective. Then any morphism $f:\mathcal{S}\to\mathcal{Y}^{\prime}$ which factors through $\pi_{F}:S\to Y$ over the generic point factors through $\pi$.
@@
Let $L^{\prime}$ be a very ample bundle on $\mathcal{Y}^{\prime}$, let $\mathcal{Y}^{j}\subset\mathcal{Y}$ be an $\mathcal{O}$-flat stratum in a flattening stratification of $\mathcal{S}\to\mathcal{Y}$, let $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ be the base-change to $\mathcal{Y}^{j}$, and consider the composition $\mathcal{S}^{j}\to\mathcal{S}\to\mathcal{Y}^{\prime}$. For any fiber $F$ of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$, $L|_{F}$ is globally generated, hence has a section which is generically nonzero on every geometric component. On the other hand, $L^{\vee}$ is trivial on every geometric fiber of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ over the generic point, hence by semicontinuity $L^{\vee}|_{F}$ has at least one nonzero section. The product is a section of $\mathcal{O}_{F}$ which is not identically zero, so since the fibers of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ are geometrically connected it must be nowhere zero. Thus, $L|_{F}$ is trivial, and so $f$ contracts every fiber of $\mathcal{S}\to\mathcal{Y}$. By the rigidity lemma, $f$ factors through $\mathcal{S}\to\mathcal{Y}$. ∎
@@
Assume $Y$ admits a local system $\mathbb{V}_{\ell}$ with faithful geometric local monodromy. Then:
@@ 2.6 Descending and extending arithmetic local systems
Let $\mathcal{O}$ be a Dedekind domain with fraction field $F$, $\mathcal{P}/\mathcal{O}$ smooth with geometrically irreducible generic fiber, and $\mathcal{Y}/\mathcal{O}$ normal finite type flat with geometrically irreducible generic fiber. Let $f:\mathcal{P}\to\mathcal{Y}$ be a proper morphism with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}f_{*}\mathcal{O}_{\mathcal{P}}$. In this section we describe how arithmetic local systems can be extended to integral models and descended along $f$.
@@
Assume $\mathcal{O}$ has perfect residue fields [^2] Over perfect fields, a scheme is geometrically reduced if and only if it is reduced.. Then after inverting finitely many primes of $\mathcal{O}$, any $\ell$-adic local system $\mathbb{V}_{\ell}$ on $\mathcal{P}$ which geometrically descends to $Y_{\bar{F}}$ descends to $\mathcal{Y}$. Moreover, the same is true for any further localization of $\mathcal{O}$.
@@
Assume $\mathcal{O}$ has perfect residue fields. Then after inverting finitely primes of $\mathcal{O}$, the following is true. For any $\ell$ and any $n$, any $\mathbb{Z}_{\ell}$-local system $\mathbb{V}_{\ell}$ on $\mathcal{P}[1/n]$ which geometrically descends to $Y_{\bar{F}}$ descends to $\mathcal{Y}[1/n]$.
@@ 2.7 Relation to Hodge theory over $\mathbb{C}$
    $\phi$ is proper if and only if and only if $V_{\mathbb{Z}}$ has faithful geometric local monodromy.
@@
    There is a finite étale cover $\pi:\tilde{P}\to P$ and a partial log smooth compactification $\tilde{P}^{\prime}$ of $\tilde{P}$ such that $\pi^{*}W_{\mathbb{Z}}$ extends to $\tilde{P}^{\prime}$ with faithful geometric local monodromy.
@@
    If the monodromy of $W_{\mathbb{Z}}$ is unipotent, then there is a partial log smooth compactification $P^{\prime}$ of $P$ such that $W_{\mathbb{Z}}$ extends to $P^{\prime}$ with failtful geometric local monodromy.
@@
For simplicity, we will assume $W_{\mathbb{Z}}$ has unipotent local monodromy. In this case, it follows from the lemma that for any period map $\phi:P\to[\mathbf{G}(\mathbb{Z})\backslash D]$, there is a partial log smooth compactification $P^{\prime}$ of $P$, a proper extension $\phi^{\prime}:P^{\prime{\mathrm{an}}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$, and an algebraic map $f:P^{\prime}\to Y$ to a normal $Y$ with $\mathcal{O}_{Y}\to f_{*}\mathcal{O}_{P^{\prime}}$ such that $\phi^{\prime}$ factors as $P^{\prime{\mathrm{an}}}\xrightarrow{f^{\mathrm{an}}}Y^{\mathrm{an}}\xrightarrow{\psi^{\mathrm{an}}}$. We call $Y$ the Stein factorization of the period map $\phi$. Observe that the local system $W_{\mathbb{Z}}$ as well as the associated filtered flat vector bundle $(W,\nabla,F^{\bullet}W)$ are pulled back from $V_{\mathbb{Z}}$ and $(V,\nabla,F^{\bullet}V)$ on $Y$.
@@
Let $Y$ be the Stein factorization of a period map $(W_{\mathbb{Z}},F^{\bullet}W)$ on $P$ with unipotent local monodromy. Let $X$ be a smooth complex algebraic variety and $U\subset X$ a dense Zariski open set. Then a morphism $\phi_{U}:U\to Y$ extends to $\phi:X\to Y$ if and only if $\pi_{U}^{*}\mathbb{W}_{\mathbb{Z}}$ extends to $X$.
@@
    If the filtered flat vector bundle $(W,\nabla,F^{\bullet}W)$ underling $(W_{\mathbb{Z}},F^{\bullet}W)$ is defined over $F$, so is the Stein factorization of the (compactified) period map $f:P^{\prime}\to Y$. Moreover, the descent $(V,\nabla,F^{\bullet}V)$ is also defined over $F$.

### v2

@@ 2.3 Natural boundary and local systems
For $\mathcal{S}/\mathcal{O}$ smooth and $\bar{\mathcal{S}}$ a log smooth compactification, both the naturality of the boundary and the admissibility condition on $\mathcal{O}$-morphisms to $\mathcal{S}$ can be checked using an $\ell$-adic local system $\mathbb{V}_{\ell}$ on $\mathcal{S}$. The key property such a local system must have is “nonextendability” over the generic fiber, meaning it does not extend to any nontrivial partial compactification—see [brunebarbe, §3], where it is referred to as “maximality”.
@@
Let $Y/\mathbb{C}$ be a normal variety with an $\ell$-adic local system $\mathbb{V}_{\ell}$. We say $\mathbb{V}_{\ell}$ is nonextendable if for any resolution $\pi:S\to Y$, finite étale $p:S^{\prime}\to S$, and partial log smooth compactification $S^{\prime}\subset\bar{S}^{\prime}$, if the local system $(p\circ\pi)^{*}\mathbb{V}_{\ell}$ extends to $\bar{S}^{\prime}$, then $S^{\prime}=\bar{S}^{\prime}$.
For $Y/F$ a normal variety with an $\ell$-adic local system $\mathbb{V}_{\ell}$, we say $\mathbb{V}_{\ell}$ is geometrically nonextenable if the base-change to $Y_{\mathbb{C}}$ is, for some (hence any) complex embedding $F\subset\mathbb{C}$. We say an $\ell$-adic local system on an integral model $\mathcal{Y}$ of $Y$ is geometrically nonextendable if the restriction to the generic fiber is.
Nonextendability restricts well:
@@
    $\mathbb{V}_{\ell}$ is nonextendable.
@@
    For any curve $C/\mathbb{C}$ and any proper morphism $f:C\to Y$, $f^{*}\mathbb{V}_{\ell}$ is nonextendable.
@@
    For any normal variety $X/\mathbb{C}$ and any proper morphism $f:X\to Y$, $f^{*}\mathbb{V}_{\ell}$ is nonextendable.
The next lemma interprets the naturality of the boundary and admissibility in terms of the local system as promised.
@@
Suppose $\mathcal{Y}/\mathcal{O}$ is a normal finite type flat $\mathcal{O}$-scheme, $\pi:\mathcal{S}\to\mathcal{Y}$ a resolution, $\bar{\mathcal{S}}/\mathcal{O}$ a log smooth compactification of $\mathcal{S}$, and $\mathbb{V}_{\ell}$ a geometrically nonextendable $\ell$-adic local system on $\mathcal{Y}$. Then
@@
    $\bar{\mathcal{S}}$ has $\pi$-natural boundary.
@@
    For $\mathcal{X}/\mathcal{O}$ regular, a morphism $f_{F}:X\to Y$ over the generic point which meets the locus of $Y$ where $\pi$ is an isomorphism is $\mathcal{O}$-admissible if and only if $f_{F}^{*}\mathbb{V}_{\ell}$ extends to $\mathcal{X}$.
@@
For 1, by Lemma 2.7, for any proper morphism $f:C\to S_{\mathbb{C}}$ from a curve, the local monodromy of $f^{*}\mathbb{V}_{\ell}$ around any boundary point is infinite.
@@ 2.4 Integral canonical models of smooth varieties
Let $\mathcal{Y}/\mathcal{O}$ a quasiprojective smooth $\mathcal{O}$-scheme. We say $\mathcal{Y}/\mathcal{O}$ is an integral canonical model if for every closed point $v$ of $\operatorname{Spec}\mathcal{O}$, the base-change $\mathcal{Y}_{v}/\mathcal{O}_{v}$ satisfies the extension property over $\mathcal{O}_{v}$.
@@
Let $\mathcal{S},\mathcal{S}^{\prime}/\mathcal{O}$ be two smooth integral canonical models and $f_{F}:S\to S^{\prime}$ a morphism over the generic point.
@@
    Suppose further that $\mathcal{S}^{\prime}$ admits a log smooth compactification and a geometrically nonextendable $\ell$-adic local system $\mathbb{V}_{\ell}$. Then $f_{F}$ extends to a morphism $f:\mathcal{S}\to\mathcal{S}^{\prime}$ if and only if $f_{F}^{*}(\mathbb{V}_{\ell})$ extends to $\mathcal{S}$.
@@ 2.5 Integral canonical models of normal varieties
Let $\mathcal{Y}/\mathcal{O}$ be a normal finite type flat $\mathcal{O}$-scheme.
@@
Let $Y/F$ be a normal variety, $\pi_{F}:S\to Y$ a resolution, and $\pi:\mathcal{S}\to\mathcal{Y}$ a model of $\pi_{F}$ defined over $\mathcal{O}$. Then $\pi$ is a uniform resolution over a (nonempty) open set of $\mathcal{O}$.
@@
Let $\mathcal{Y}/\mathcal{O}$ a normal quasiprojective flat $\mathcal{O}$-scheme. We say $\mathcal{Y}/\mathcal{O}$ is an integral canonical model if:
@@
    Suppose further that $\mathcal{Y}^{\prime}$ admits a uniform resolution $\pi^{\prime}:\mathcal{S}^{\prime}\to\mathcal{Y}^{\prime}$, a log smooth compactification of $\mathcal{S}^{\prime}$, and that $\mathcal{Y}^{\prime}$ has a geometrically nonextendable $\ell$-adic local system $\mathbb{V}_{\ell}$. Then $f_{F}$ extends to a morphism $f:\mathcal{Y}\to\mathcal{Y}^{\prime}$ if and only if $f_{F}^{*}(\mathbb{V}_{\ell})$ extends to $\mathcal{Y}$.
@@
Let $\mathcal{Y},\mathcal{Y}^{\prime}/\mathcal{O}$ be normal finite type flat $\mathcal{O}$-schemes, and $\pi:\mathcal{S}\to\mathcal{Y}$ a uniform resolution. Assume $\mathcal{Y}^{\prime}/\mathcal{O}$ is quasiprojective. Then any morphism $f:\mathcal{S}\to\mathcal{Y}^{\prime}$ which factors through $\pi_{F}:S\to Y$ over the generic point factors through $\pi$.
@@
Let $L^{\prime}$ be a very ample bundle on $\mathcal{Y}^{\prime}$, let $\mathcal{Y}^{j}\subset\mathcal{Y}$ be an $\mathcal{O}$-flat stratum in a flattening stratification of $\mathcal{S}\to\mathcal{Y}$, let $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ be the base-change to $\mathcal{Y}^{j}$, and consider the composition $\mathcal{S}^{j}\to\mathcal{S}\to\mathcal{Y}^{\prime}$. For any fiber $F$ of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$, $L|_{F}$ is globally generated, hence has a section which is generically nonzero on every geometric component. On the other hand, $L^{\vee}$ is trivial on every geometric fiber of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ over the generic point, hence by semicontinuity $L^{\vee}|_{F}$ has at least one nonzero section. The product is a section of $\mathcal{O}_{F}$ which is not identically zero, so since the fibers of $\mathcal{S}^{j}\to\mathcal{Y}^{j}$ are geometrically connected it must be nowhere zero. Thus, $L|_{F}$ is trivial, and so $f$ contracts every fiber of $\mathcal{S}\to\mathcal{Y}$. By the Lemma 2.17, we conclude that $f$ factors through $\mathcal{S}\to\mathcal{Y}$. ∎
@@
Assume $Y$ admits a geometrically nonextendable local system $\mathbb{V}_{\ell}$. Then:
@@ 2.6 Descending and extending arithmetic local systems
Let $\mathcal{P}/\mathcal{O}$ be smooth with geometrically irreducible generic fiber, and $\mathcal{Y}/\mathcal{O}$ normal finite type flat with geometrically irreducible generic fiber. Let $f:\mathcal{P}\to\mathcal{Y}$ be a proper morphism of $\mathcal{O}$-schemes with $\mathcal{O}_{\mathcal{Y}}\xrightarrow{\cong}f_{*}\mathcal{O}_{\mathcal{P}}$. In this section we describe how arithmetic local systems can be extended to integral models and descended along $f$.
@@
After inverting finitely many primes of $\mathcal{O}$, any $\ell$-adic local system $\mathbb{V}_{\ell}$ on $\mathcal{P}$ which geometrically descends to $Y_{\bar{F}}$ descends to $\mathcal{Y}$. Moreover, the same is true for any further localization of $\mathcal{O}$.
@@
After inverting finitely primes of $\mathcal{O}$, the following is true. For any $\ell$ and any $n$, any $\mathbb{Z}_{\ell}$-local system $\mathbb{V}_{\ell}$ on $\mathcal{P}[1/n]$ which geometrically descends to $Y_{\bar{F}}$ descends to $\mathcal{Y}[1/n]$.
@@ 2.7 Relation to Hodge theory over $\mathbb{C}$
    $\phi$ is proper if and only if and only if $V_{\mathbb{Z}}$ is nonextendable [^3] We use the same definition for a $\mathbb{Z}$-local system.
@@
    There is a finite étale cover $\pi:\tilde{P}\to P$ and a partial log smooth compactification $\tilde{P}^{\prime}$ of $\tilde{P}$ such that $\pi^{*}W_{\mathbb{Z}}$ extends to $\tilde{P}^{\prime}$ where it is nonextendable.
@@
    If the local monodromy of $W_{\mathbb{Z}}$ is unipotent, then there is a partial log smooth compactification $P^{\prime}$ of $P$ such that $W_{\mathbb{Z}}$ extends to $P^{\prime}$ where it is nonextendable.
@@
For simplicity, we will assume $W_{\mathbb{Z}}$ has torsion-free monodromy. In this case, it follows from the lemma that for any period map $\Phi:P^{\mathrm{an}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$, there is a partial log smooth compactification $P^{\prime}$ of $P$, a proper extension $\Phi^{\prime}:P^{\prime{\mathrm{an}}}\to[\mathbf{G}(\mathbb{Z})\backslash D]$ which factors through a finite inertia-free cover $\Gamma\backslash D$, and a proper algebraic map $h^{\prime}:P^{\prime}\to Y$ to a normal variety $Y$ with $\mathcal{O}_{Y}\xrightarrow{\cong}h^{\prime}_{*}\mathcal{O}_{P^{\prime}}$ such that $\Phi^{\prime}$ factors as $P^{\prime{\mathrm{an}}}\xrightarrow{h^{\prime{\mathrm{an}}}}Y^{\mathrm{an}}\xrightarrow{\psi^{\mathrm{an}}}[\mathbf{G}\backslash D]$. We call $Y$ the Stein factorization of the period map $\Phi$. Observe that the local system $W_{\mathbb{Z}}$ as well as the associated filtered flat vector bundle $(W,\nabla,F^{\bullet}W)$ are pulled back from $V_{\mathbb{Z}}$ and $(V,\nabla,F^{\bullet}V)$ on $Y$.
@@
Let $Y$ be the Stein factorization of a period map $(W_{\mathbb{Z}},F^{\bullet}W)$ on $P$ with torsion-free monodromy. Let $X$ be a smooth complex algebraic variety and $U\subset X$ a dense Zariski open set. Then a morphism $g_{U}:U\to Y$ extends to $g:X\to Y$ if and only if $g_{U}^{*}\mathbb{W}_{\mathbb{Z}}$ extends to $X$.
Definition 2.27.
For any scheme $Y$ and a vector bundle $(V,F^{\bullet}V)$ on $Y$ with a locally split decreasing filtration, we define the Griffiths bundle of $(V,F^{\bullet}V)$ to be $\otimes_{i\in\mathbb{Z}}\det F^{i}V$.
The main result of [bbt] is the following:
Theorem 2.28 ([bbt]).
If $Y$ is the Stein factorization of a period map as above, then the Griffiths bundle on $Y$ is ample.
@@
    If the filtered flat vector bundle $(W,\nabla,F^{\bullet}W)$ underling $(W_{\mathbb{Z}},F^{\bullet}W)$ is defined over $F$, so is the Stein factorization of the (compactified) period map $h^{\prime}:P^{\prime}\to Y$. Moreover, the descent $(V,\nabla,F^{\bullet}V)$ is also defined over $F$.